OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y <= -2x + 8? (0, -4) (5, -1) (7, 0) (9, 1)

OpenStudy (ash2326):

@kaddahannas We have \[y\le -2x+8\] we need to check each of the options, I'll check the last one for you. (9, 1) \[1\le -2\times 9+8\] \[1\le -18+8\] \[1\le -10\] This is not true \[1\cancel {\le} -10\] Would you check other options?

OpenStudy (anonymous):

is it the second one?

OpenStudy (ash2326):

no wait

OpenStudy (ash2326):

it's \(-1\le -2\) which is not true. Please check again

OpenStudy (anonymous):

oh wait its the first one correct?

OpenStudy (ash2326):

yeah \[-4\le 8\ \] It's correct.:)